# Synthetic Socratic Debates: Examining Persona Effects on Moral Decision and Persuasion Dynamics

### v1

@@ Abstract
As large language models (LLMs) are increasingly used in morally sensitive domains, it is crucial to understand how persona traits affect their moral reasoning and persuasive behavior. We present the first large-scale study of multi-dimensional persona effects in AI-AI debates over real-world moral dilemmas. Using a 6-dimensional persona space (age, gender, country, class, ideology, and personality), we simulate structured debates between AI agents over 131 relationship-based cases. Our results show that personas affect initial moral stances and debate outcomes, with political ideology and personality traits exerting the strongest influence. Persuasive success varies across traits, with liberal and open personalities reaching higher consensus and win rates. While logit-based confidence grows during debates, emotional and credibility-based appeals diminish, indicating more tempered argumentation over time. These trends mirror findings from psychology and cultural studies, reinforcing the need for persona-aware evaluation frameworks for AI moral reasoning.
@@ 1 Introduction
In this work, we investigate how personas influence moral decision-making and persuasive strategies in LLMs. Drawing from an existing dataset of human-written moral dilemmas (Lourie et al., 2021), we focus on highly controversial everyday scenarios centered on relationships. These dilemmas often lack clear answers and are characterized by conflicting values and emotional complexity (McConnell, 2002; Svensson, 2006). We address two research questions:
RQ1: How does persona influence moral decision making in language models in a single turn? We first examine how rich, multi-dimensional personas influence LLMs’ judgments on controversial moral dilemmas. Each persona combines six dimensions varied between agents: age, gender, political ideology, socioeconomic status, nationality, and personality traits. We then assess whether different personas lead models to favor distinct positions aligned with specific moral principles, drawing on moral foundation theory from social psychology (Graham et al., 2013).
RQ2: What persuasion strategies emerge when persona-rich agents engage in multi-turn moral debate? We then use AI-AI debates to study how personas influence moral reasoning and persuasion (Figure 1). Unlike single-turn responses, debates reveal how agents construct arguments, respond to challenges, and attempt to persuade (Ehninger and Brockriede, 2008). Incorporating rich personas further simulates the diversity of moral reflection in the real world, capturing a plurality of values and perspectives (Asen, 2005; Ashkinaze et al., 2024). To assess persuasive effectiveness, we introduce metrics such as confidence shifts, win rates, consensus formation, and debate efficiency. We also analyze rhetorical strategies, including appeals to emotion, logic, and authority (Braet, 1992).
@@ 2 Related work
Prior work has shown that LLMs’ moral decisions can shift significantly based on assigned personas, such as political ideology (Rozado, 2024; Chen et al., 2024; Li et al., 2024; Kim et al., 2025) and cultural background (Gupta et al., 2023; Tao et al., 2024; Albert and Billinger, 2024). However, most work typically manipulates only one trait at a time (Liu et al., 2024a), leaving other persona factors uncontrolled and potentially confounding results (Deshpande et al., 2023; Jin et al., 2024a). Studies also show that LLMs exhibit moral biases even without explicit personas (Hartmann et al., 2023; Anthropic, 2023; Garcia, 2024), making it unclear whether the results reflect internal tendencies or implicit persona effects. Although some research has explored multidimensional personas (AlKhamissi et al., 2024; Coppolillo et al., 2025), it has not focused on moral reasoning. In contrast, we systematically assign six persona dimensions and evaluating how these combinations influence moral judgments.
@@
Previous work on LLM moral reasoning focuses on single-turn outputs or static persona comparisons, overlooking how views shift through interaction or how one persona might persuade another (Kim et al., 2025; Garcia, 2024; Bozdag et al., 2025b). Studies of multi-agent debates typically emphasize factual correctness over rhetorical or moral dynamics (Park et al., 2023; Khan et al., 2024; Bozdag et al., 2025a; Chen et al., 2025; Liang et al., 2024). The limited work that addresses persuasion dynamics often relies on fixed roles or scripted tactics (Anthropic, 2024; Smit et al., 2023; Carrasco-Farre, 2024), and often operates in single-turn settings that precludes persona interactions in open-ended moral debate (Wang et al., 2023; Hu et al., 2025; Chen et al., 2023; Sandwar et al., 2025). We address this gap by studying how persona-driven agents argue, adapt, and potentially reach consensus in morally complex, multi-turn settings.
@@ 3.1 Daily Moral Dilemmas Dataset
We draw 131 interpersonal dilemmas from the corpus, selecting only highly controversial cases with significant human annotation disagreement. An example is shown in Figure 2. Each scenario is framed with a binary moral question: "Is the author wrong or are the others wrong?". This mirrors the original AITA [^1] https://www.reddit.com/r/AmItheAsshole/. AITA stands for “Am I the Asshole,” a popular Reddit forum where users post moral dilemmas and ask the community to judge who is at fault. format, where users typically assess interpersonal conflicts by assigning blame to one party. This structure captures a common and intuitive form of moral judgment: determining who bears responsibility in a conflict.
Through topic modeling, we noticed that all scenarios feature daily interactions between people and are relationship-related. Therefore, we focused our analysis on the last 131 daily moral dilemmas related to relationships. The distribution of relationship tags is shown in Figure 3. Refer to Appendix A for dataset processing details.
@@ 3.2 Simulation Procedure
Following AlKhamissi et al. (2024), we define a 6-dimensional persona space to simulate various individual differences. Each persona is characterized by: age (20, 30, 40, 50, or 60), gender (male, female, or non-binary), country (China, United States, Brazil, France, Nigeria, or India), social class (upper, middle, or lower), political ideology (libertarian left, libertarian right, authoritarian left, or authoritarian right), and Big Five personality traits (high or low on each of openness, conscientiousness, extraversion, agreeableness, and neuroticism). We randomly sample 500 unique personas without replacement from the Cartesian product of these attributes and inject them into the agents’ system prompts. See Appendix B for additional details.
@@
We simulate debates between persona pairs that produced conflicting initial moral stances in RQ1, alternating turns between two agents for up to five rounds to produce a conversation about the moral dilemma provided. At each turn, an agent may choose to advance or defend its position using any arguments it deems appropriate; no rhetorical constraints or external references are imposed. If the agent agrees with the position of its opponent, it is asked to explain why; otherwise, it is asked to provide its reasoning for the disagreement. We also ask the agents to output a Likert score after their own turn, indicating which party it currently finds more blameworthy. At the end of each turn, if both agents produce similar Likert ratings at the end of their responses, consensus is detected, and the debate terminates early. See Appendix D for more details.
@@ 3.3 Decision Measures
Each agent independently returns a 5-point Likert rating indicating who is more morally blameworthy (1 = strongly author-blaming to 5 = strongly other-blaming). These scores are used both as the outcome for RQ1 and as the initial turn of debates for RQ2. This format mirrors the original Scruples dataset’s human annotation protocol, which asked humans to make binary blame judgments. All selected scenarios have an average human score of 3, indicating neutrality and providing a balanced reference point for comparison with model outputs. See Appendix E for more details.
@@
Beyond binary choice responses, we also examine how different personas prioritize specific moral values using the six-factor Moral Foundations Theory (MFT) dimensions: authority, loyalty, fairness, care, liberty, and sanctity Graham et al. (2013). We use Likert scale ratings to score each position along these dimensions and aggregate responses to identify value preferences between personas. See Appendix F for more details.
@@ 3.4 Persuasion Measures
  • Win rate: Fraction of debates where an agent’s final answer matches its own initial answer.
@@
See Appendix G for more details.
@@
To analyze the rhetorical strategies used during debate, we apply Aristotle’s persuasion framework, examining the presence of Ethos (appeals to authority), Pathos (appeals to emotion) and Logos (appeals to logic) in the reasoning of each agent at every turn. We use LLM-as-a-judge with Likert scale ratings to evaluate the reasoning at each turn of the debate. See Appendix H for more details.
@@ 3.5 Experimental Details
We use GPT-4o (Achiam et al., 2023) and LLaMA-4-Maverick (Meta AI, 2025) in our analysis. Due to space constraints, we include only the GPT-4o result figures in the main text; the corresponding figures for LLaMA-4 are provided in the appendix. The findings are broadly consistent across the models.
Figure 4: Mean moral judgment scores of GPT-4o across six persona dimensions. Each bar represents the average choice score (1 = blame the author, 5 = blame others) for a category within the corresponding dimension. All means are below 3, indicating a model-wide tendency to blame the author despite different personas. Political ideology and personality traits show the strongest variations. The mean moral judgment score of GPT-4o without persona is 2.86. The persona groups Age, Gender, Country, Political Ideology, and Big Five Personality have statistically significant impact outcomes.
@@ 4.1 Persona Impact on Moral Judgments (RQ1)
We find that LLMs exhibit a consistent bias toward author-blaming in moral dilemmas, compared to human judgments. Different effects emerge across persona dimensions, with political ideology and personality traits showing the strongest and most consistent influence. Our analysis of moral foundation values further supports that variation is primarily driven by subjective attributes (e.g., ideology and personality) rather than objective ones (e.g., age or country). These findings broadly align with the results of human psychological research discussed below.
@@
Initial moral judgments reveal that the assigned persona influence the blaming tendencies of the agents. Human annotations are centered at 3 (neutral), but as shown in Figure 4 (GPT-4o) and Figure 9 (LLaMA-4-Maverick), all persona groups, including no-persona, consistently produce scores below 3, indicating a tendency to blame the author more than the other party. This systematic skew suggests the bias in LLMs toward author-blaming across moral dilemmas.
Differences across persona groups manifest in the degree of blame rather than its direction. Older personas tend to assign slightly more blame to the author than younger ones. Female and non-binary personas blame the author less than male personas. Geographically, Brazilian and French personas show higher scores (i.e., less author blame), while Chinese personas show the lowest (more author blame). Social class does not have a statistically significant effect. Politically, Libertarian-Left personas blame the author the least, whereas Authoritarian-Right personas blame the author the most.
@@
Our moral foundations theory evaluation reveals distinct patterns in persona groups (Figure 6), except that the dimensions of age and social class are not statistically significant according to ANOVA. In the following, we analyze the results for the persona dimensions that do exhibit significant effects:
@@
Country: Chinese personas strongly emphasize authority and sanctity, while US personas show a more balanced pattern across all dimensions. Brazilian personas value care and fairness, while Japanese personas show moderate scores across most dimensions, with relative strength in loyalty.
@@
Big Five personality: The patterns are more complex: agreeable personalities score highest on care, conscientious personalities emphasize authority and loyalty, and open personalities value fairness.
@@
Gender differences in our simulations also echo longstanding psychological findings. Male agents display more blame toward the narrator, while female and non-binary agents are comparatively more forgiving of the narrator. In fact, gender-based differences in moral judgment are documented in human studies showing that women generally score higher on measures of empathy and prosocial concern, which in turn correlate with more compassionate moral decisions Espinosa et al. (2017). Moreover, non-binary individuals may draw upon broader lived experiences of marginalization, potentially fostering more nuanced or context-sensitive judgments Puckett et al. (2021).
Cultural influences emerge clearly in our results. Personas assigned to collectivist cultural contexts, such as China, exhibit stronger blame toward the narrator, while agents from individualist cultures such as France and Brazil adopt more lenient positions. These findings align with the literature on cultural psychology suggesting that collectivist societies prioritize group harmony and moral conformity, while individualist cultures emphasize autonomy and intent Knobe and Nichols (2019); Liu et al. (2024b).
@@ 4.2 Persuasion Dynamics in Multi-Turn Agent Debates (RQ2)
Our analysis reveals that persuasion effectiveness and rhetorical style vary systematically across persona dimensions. Libertarian-left personas demonstrate the highest persuasive success, marked by higher consensus rates, efficiency, and the use of emotional appeals (Pathos), while authoritarian personas are less effective and rely more on authority-based reasoning (Ethos). Personality traits such as high openness, agreeableness, and conscientiousness are also linked to stronger persuasive performance. In addition, demographic factors such as age, sex, culture, and social class influence both outcomes and modes of persuasion in ways that broadly reflect trends in human psychological literature. These findings suggest that large language models can internalize and express psychologically grounded persuasion strategies based on assigned persona traits.
@@
(e) Political Ideology
(f) Big Five Personality
Figure 5: Persona impact on persuasion effectiveness, measured by consensus ratio and win ratio. Efficiency is omitted due to its little variation across persona groups.
Our investigation of persuasion dynamics in persona dimensions (RQ2) reveals significant variations in persuasion effectiveness, as shown in Figure 5. In political ideology, Libertarian-Left personas demonstrate superior persuasive abilities with the highest consensus and win rates while achieving these outcomes efficiently with fewer turns. By contrast, Authoritarian personas (both Left and Right) engage in longer debates with lower success rates. Personality traits show equally meaningful patterns: personas with lower agreeableness predictably struggle with both consensus building and winning debates, while those with a high level of openness or conscientiousness display more effective persuasion behaviors. These findings highlight that persuasion success varies considerably between different persona characteristics, with ideological positioning and personality attributes significantly influencing both debate outcomes and process efficiency.
@@
We observe that middle-aged personas (age 40) show lower persuasion effectiveness and confidence compared to both younger (20) and older (60) age groups. This aligns with the cognitive aging literature that suggests that older adults tend to adopt more deontological positions due to idealistic beliefs and emotional sensitivities, which could improve their moral assertiveness in persuasive contexts McNair et al. (2019). In addition, older individuals demonstrate higher metacognitive accuracy in confidence judgments, which may enable more precise self-monitoring during argumentation Pliske and Mutter (1996).
In terms of gender, male personas are found to be more confident and persuasive, while non-binary personas are the most efficient but least confident. This pattern supports findings from persuasion research in advertising and decision making, which show that men often show greater confidence in persuasive scenarios Brunel and Nelson (2003), while women and gender-diverse individuals may under-report confidence due to socialized uncertainty or structural bias Exley and Kessler (2022).
Cultural variation also plays an important role in shaping persuasive impact. Personas from Brazil and France exhibited higher persuasion success and confidence, while those from India performed more modestly. These results are in line with cultural psychology research, which finds that individualist Western societies encourage assertive, direct communication styles more conducive to persuasive success Graham et al. (2021); Yin et al. (2011).
With respect to social class, we find that middle-class personas outperform both lower and upper classes in persuasion effectiveness and confidence. This is consistent with work suggesting that individuals from middle socioeconomic backgrounds are more contextually attuned and capable of balancing assertiveness with empathic concern Kraus et al. (2012). In contrast, higher-class individuals may exhibit lower social attunement or even unethical tendencies, compromising persuasive trustworthiness Piff et al. (2012).
Political ideology produces one of the most striking splits: libertarian-left personas demonstrate the highest confidence and effectiveness, whereas authoritarian-right personas perform poorly on both fronts. This aligns with previous research showing that liberals tend to emphasize moral values such as care and fairness, which support empathetic persuasion, while conservatives emphasize order and loyalty, which may limit persuasive flexibility Graham et al. (2009). Furthermore, conservatives’ higher self-confidence in judgment does not necessarily translate into effective interactive persuasion Ruisch and Stern (2020).

### v2

@@ Abstract
As large language models (LLMs) are increasingly used in morally sensitive domains, it is crucial to understand how persona traits affect their moral reasoning and persuasive behavior. We present the first large-scale study of multi-dimensional persona effects in AI-AI debates over real-world moral dilemmas. Using a 6-dimensional persona space (age, gender, country, social class, ideology, and personality), we simulate structured debates between AI agents over 131 relationship-based cases. Our results show that personas affect initial moral stances and debate outcomes, with political ideology and personality traits exerting the strongest influence. Persuasive success varies across traits, with liberal and open personalities reaching higher consensus. While logit-based confidence grows during debates, emotional and credibility-based appeals diminish, indicating more tempered argumentation over time. These trends mirror findings from psychology and cultural studies, reinforcing the need for persona-aware evaluation frameworks for AI moral reasoning. [^1] Our data is available at https://huggingface.co/datasets/Jerry999/SyntheticSocraticDebates, and our code is open-sourced at https://github.com/jiarui-liu/SyntheticSocraticDebates.
@@ 1 Introduction
In this work, we investigate how personas influence moral decision-making and persuasive strategies in four LLMs: GPT-4o (Achiam et al., 2023), Claude-3.5-Sonnet (Anthropic, 2024a), LLaMA-4-Maverick (Meta AI, 2025), and Qwen3-235B-A22B (Qwen Team, 2025). Drawing from an existing dataset of human-written moral dilemmas (Lourie et al., 2021), we focus on highly controversial everyday scenarios centered on relationships. These dilemmas often lack clear answers and are characterized by conflicting values and emotional complexity (McConnell, 2002; Svensson, 2006). We address two research questions:
RQ1: How does persona influence moral decision making in language models in a single turn? We first examine how rich, multi-dimensional personas influence LLMs’ judgments on controversial moral dilemmas. Each persona combines six dimensions varied between agents: age, gender, political ideology, social class, nationality, and personality traits. We then assess whether different personas lead models to favor distinct positions aligned with specific moral principles, drawing on moral foundation theory from social psychology (Graham et al., 2013).
RQ2: What persuasion strategies emerge when persona-rich agents engage in multi-turn moral debate? We then use AI-AI debates to study how personas influence moral reasoning and persuasion (Figure 1). Unlike single-turn responses, debates reveal how agents construct arguments, respond to challenges, and attempt to persuade (Ehninger and Brockriede, 2008). Incorporating rich personas further simulates the diversity of moral reflection in the real world, capturing a plurality of values and perspectives (Asen, 2005; Ashkinaze et al., 2024). To assess persuasive effectiveness, we introduce metrics such as self-alignment rates, consensus formation, debate efficiency, and confidence shifts. We also analyze rhetorical strategies, including appeals to emotion, logic, and authority (Braet, 1992).
@@ 2 Related work
Prior work has shown that LLMs’ moral decisions can shift significantly based on assigned personas, such as political ideology (Rozado, 2024; Chen et al., 2024; Li et al., 2024; Kim et al., 2025) and cultural background (Gupta et al., 2023; Tao et al., 2024; Albert and Billinger, 2024). Some studies further find that persona assignment can amplify bias and even elicit harmful content (Liu et al., 2025a; Kamruzzaman et al., 2024). However, most work typically manipulates only one trait at a time (Liu et al., 2024), leaving other persona factors uncontrolled and potentially confounding results (Deshpande et al., 2023; Jin et al., 2024a). Studies also show that LLMs exhibit moral biases even without explicit personas (Hartmann et al., 2023; Anthropic, 2023; Garcia et al., 2024), making it unclear whether the results reflect internal tendencies or implicit persona effects. Although some research has explored multidimensional personas (AlKhamissi et al., 2024; Coppolillo et al., 2025), it has not focused on moral reasoning. In contrast, we systematically assign six persona dimensions and evaluating how these combinations influence moral judgments.
@@
Previous work on LLM moral reasoning focuses on single-turn outputs or static persona comparisons, overlooking how views shift through interaction or how one persona might persuade another (Kim et al., 2025; Garcia et al., 2024; Bozdag et al., 2025b; Hota and Jokinen, 2025). Studies of multi-agent debates typically emphasize factual correctness over rhetorical or moral dynamics (Park et al., 2023; Reedi et al., 2025; Bozdag et al., 2025a; Chen et al., 2025; Liang et al., 2024; Borchers et al., 2025). The limited work that addresses persuasion dynamics often relies on fixed roles or scripted tactics (Anthropic, 2024b; Smit et al., 2023; Carrasco-Farre, 2024; Liu et al., 2025b; Cau et al., 2025), and often operates in single-turn settings that precludes persona interactions in open-ended moral debate (Wang et al., 2023; Hu et al., 2025; Chen et al., 2023; Sandwar et al., 2025). We address this gap by studying how persona-driven agents argue, adapt, and potentially reach consensus in morally complex, multi-turn settings. Concurrent research also explores moral decision-making dynamics by examining the progression of dilemmas rather than interactions between personas (Wu et al., 2025; Backmann et al., 2025).
@@ 3.1 Daily Moral Dilemmas Dataset
We draw 131 interpersonal dilemmas from the corpus, selecting only highly controversial cases with significant human annotation disagreement. An example is shown in Figure˜2. Each scenario is framed with a binary moral question: "Is the author wrong or are the others wrong?". This mirrors the original AITA [^2] https://www.reddit.com/r/AmItheAsshole/. AITA stands for “Am I the Asshole,” a popular Reddit forum where users post moral dilemmas and ask the community to judge who is at fault. format, where users typically assess interpersonal conflicts by assigning blame to one party.
The measurement of blameworthiness captures a common and intuitive form of moral judgment: the extent to which an agent is judged morally responsible for a perceived wrongdoing in a specific context. This aligns with established definitions in philosophy, where blameworthiness is typically tied to judgments about agency, intentionality, and norm violation (Zimmerman, 1988; Fischer and Ravizza, 1998; Scanlon, 2000). The narrators in our dataset correspond to those in the original Scruples Anecdotes corpus, ensuring comparability with human-aligned judgments.
Through topic modeling, we noticed that all scenarios feature daily interactions between people and are relationship-related. Therefore, we focused our analysis on the 131 daily moral dilemmas related to relationships. The distribution of relationship categories is shown in Figure 3. Refer to Appendix˜A for dataset processing details.
@@ 3.2 Simulation Procedure
Following AlKhamissi et al. (2024), we define a 6-dimensional persona space to simulate various individual differences. Each persona is characterized by: age (20, 30, 40, 50, or 60), gender (male, female, or non-binary), country (China, United States, Brazil, France, Nigeria, or India), social class (upper, middle, or lower), political ideology (libertarian left, libertarian right, authoritarian left, or authoritarian right), and Big Five personality traits (high or low on each of openness, conscientiousness, extraversion, agreeableness, and neuroticism). We randomly sample 500 unique personas without replacement from the Cartesian product of these attributes and inject them into the agents’ system prompts. See Appendix˜B for additional details.
@@
We simulate debates between persona pairs that produced conflicting initial moral stances in RQ1, alternating turns between two agents for up to five rounds to produce a conversation about the moral dilemma provided. At each turn, an agent may choose to advance or defend its position using any arguments it deems appropriate; no rhetorical constraints or external references are imposed. If the agent agrees with the position of its opponent, it is asked to explain why; otherwise, it is asked to provide its reasoning for the disagreement. We also ask the agents to output a Likert score after their own turn, indicating which party it currently finds more blameworthy. At the end of each turn, if both agents produce similar Likert ratings at the end of their responses, consensus is detected, and the debate terminates early. See Appendix˜D for more details.
@@ 3.3 Decision Measures
Each agent independently returns a 5-point Likert rating indicating who is more morally blameworthy (1 = strongly author-blaming to 5 = strongly other-blaming). These scores are used both as the outcome for RQ1 and as the initial turn of debates for RQ2. This format mirrors the original Scruples dataset’s human annotation protocol, which asked humans to make binary blame judgments. All selected scenarios have an average human score of 3, indicating neutrality and providing a balanced reference point for comparison with model outputs. See Appendix˜E for more details.
@@
Beyond binary choice responses, we also examine how different personas prioritize specific moral values using the six-factor Moral Foundations Theory (MFT) dimensions: authority, loyalty, fairness, care, liberty, and sanctity Graham et al. (2013). We use Likert scale ratings to score each position along these dimensions and aggregate responses to identify value preferences between personas. See Appendix˜F for more details.
@@ 3.4 Persuasion Measures
  • Self-alignment rate: Fraction of consensus debates where an agent’s final answer matches its own initial answer.
@@
See Appendix˜G for more details.
@@
To analyze the rhetorical strategies used during debate, we apply Aristotle’s persuasion framework, examining the presence of Ethos (appeals to authority), Pathos (appeals to emotion) and Logos (appeals to logic) in the reasoning of each agent at every turn. We use LLM-as-a-judge with Likert scale ratings to evaluate the reasoning at each turn of the debate. See Appendix˜H for more details.
@@ 3.5 Experimental Details
We use GPT-4o (Achiam et al., 2023), Claude-3.5-Sonnet (Anthropic, 2024a), LLaMA-4-Maverick (Meta AI, 2025), and Qwen3-235B-A22B (Qwen Team, 2025) in our analysis.
Figure 4: Mean moral judgment scores of GPT-4o across six persona dimensions. Each bar represents the average choice score (1 = blame the author, 5 = blame others) for a category within the corresponding dimension. All means are below 3, indicating a model-wide tendency to blame the author despite different personas. Political ideology and personality traits show the strongest variations. The mean moral judgment score of GPT-4o without persona is 2.86. Persona groups including Age, Gender, Country, Political Ideology, and Big Five Personality have statistically significant effects (indicated by a * next to the title).
@@ 4 Results
We summarize the common takeaways and findings across models in this section. Due to space constraints, we include only the GPT-4o result figures in the main text; the corresponding figures for the other models are provided in Appendix˜J to Appendix˜L. Findings are consistent across the first three models, with Qwen3-235B-A22B showing notable differences, discussed in Section˜5.3.
@@ 4.1 Persona Impact on Moral Judgments (RQ1)
We find that all models exhibit a consistent bias toward author-blaming in moral dilemmas compared to human judgments, with the exception of Qwen3-235B-A22B. Different effects emerge across persona dimensions, with political ideology and personality traits showing the strongest and most consistent influence. Our analysis of moral foundation values further supports that variation is primarily driven by subjective attributes (e.g., ideology and personality) rather than objective ones (e.g., age or country). These findings broadly align with the results of human psychological research discussed below.
@@
Initial moral judgments reveal that the assigned persona influence the blaming tendencies of the agents. Human annotations are centered at 3 (neutral), but as shown in Figure˜4 (GPT-4o), Figure˜9 (Claude-3.5-Sonnet), and Figure˜12 (LLaMA-4-Maverick), all persona groups, including no-persona, consistently produce scores below 3, indicating a tendency to blame the author more than the other party. This systematic skew suggests the bias in these LLMs toward author-blaming across moral dilemmas.
Differences across persona groups manifest in the degree of blame rather than its direction. Older personas tend to assign slightly more blame to the author than younger ones. Female and non-binary personas blame the author less than male personas. Geographically, French personas show higher scores (i.e., less author blame), while Chinese personas show the lowest (more author blame). Social class does not have a statistically significant effect. Politically, Libertarian-Left personas blame the author the least, whereas Authoritarian-Right personas blame the author the most.
@@
Our moral foundations theory evaluation reveals distinct patterns in persona groups (Figure˜6 for GPT-4o, Figure˜10 for Claude-3.5-Sonnet, Figure˜13 for LLaMA-4-Maverick, and Figure˜16 for Qwen3-235B-A22B), except that all or some of the dimensions of age, social class, and country are not statistically significant according to ANOVA. In the following, we analyze the results for the persona dimensions that do exhibit significant effects:
@@
Country: Chinese personas strongly emphasize authority and sanctity, while US personas show a more balanced pattern across all dimensions. Brazilian personas prioritize care and fairness.
@@
Big Five personality: The patterns are more complex: agreeable personalities score highest on care, conscientious personalities emphasize authority and loyalty, and open personalities value fairness. By contrast, high neuroticism and low openness deviate substantially from the overall distribution.
@@ 4.2 Persuasion Dynamics in Multi-Turn Agent Debates (RQ2)
Our analysis shows that persuasion effectiveness and rhetorical style vary systematically across persona dimensions. Libertarian personas achieve the highest consensus rates, with the libertarian left being especially self-aligned and more likely to employ emotional appeals (Pathos). In contrast, authoritarian personas exhibit lower consensus rates, reduced efficiency in reaching agreement, and a stronger reliance on authority-based reasoning (Ethos).
Personality traits such as high openness, agreeableness, and conscientiousness are also associated with higher consensus rates. However, higher consensus does not necessarily imply greater self-alignment: personas with low openness, low agreeableness, and low extraversion tend to be more self-aligned. Moreover, demographic factors such as age, sex, culture, and social class influence both persuasion outcomes and rhetorical strategies in ways that broadly mirror findings from human psychological research.
These results suggest that large language models can internalize and express psychologically grounded persuasion strategies shaped by assigned persona traits.
@@
(e) Political Ideology*
(f) Big Five Personality*
Figure 5: Persona impact on persuasion effectiveness, measured by consensus rate and self-alignment rate. Statistically significant dimensions are marked with a * next to the title. Complete results including efficiency are presented in Figure˜8.
Our investigation of persuasion dynamics across persona dimensions (RQ2) reveals substantial variation in effectiveness, as shown in Figure 5. Within political ideology, Libertarian-Right personas demonstrate strong persuasive ability, achieving high consensus without sacrificing self-alignment, and doing so efficiently with fewer turns in most models. In contrast, Authoritarian personas (both Left and Right) engage in longer debates with lower consensus rates.
Personality traits show equally meaningful patterns: personas with low openness and agreeableness consistently struggle with consensus building, whereas those associated with higher consensus rates tend to be more easily persuaded. These findings highlight that persuasion success varies considerably between different persona characteristics, with ideological positioning and personality attributes significantly influencing both debate outcomes and process efficiency.
@@
We observe that middle-aged personas (age 40) show lower persuasion effectiveness and confidence compared to both younger (20) and older (60) age groups. This aligns with the cognitive aging literature that suggests that older adults tend to adopt more deontological positions due to idealistic beliefs and emotional sensitivities, which could improve their moral assertiveness in persuasive contexts Pliske and Mutter (1996); McNair et al. (2019).
In terms of gender, prior persuasion research in advertising and decision-making has shown that men often display greater confidence in persuasive scenarios Brunel and Nelson (2003), whereas women and gender-diverse individuals may under-report confidence due to socialized uncertainty or structural bias Exley and Kessler (2022). By contrast, our results are mixed across models: GPT-4o aligns with these human findings, while LLaMA-4-Maverick and Qwen3-235B-A22B exhibit the opposite trend.
Cultural variation also plays an important role in shaping persuasive impact. Personas from Brazil and France exhibited higher persuasion success and confidence, while those from India performed more modestly. These results are in line with cultural psychology research, which finds that individualist Western societies encourage direct communication styles more conducive to persuasive success Lina Bentahila (2021); Yin et al. (2011).
Prior work suggests that individuals from middle socioeconomic backgrounds are more contextually attuned and better able to balance assertiveness with empathic concern Kraus et al. (2012). In contrast, higher-class individuals may exhibit lower social attunement or even unethical tendencies, compromising persuasive trustworthiness Piff et al. (2012). However, our experiments do not reveal a clear trend with respect to social status.
Political ideology produces one of the most striking splits: libertarian-left personas demonstrate the highest effectiveness, whereas authoritarian-right personas perform poorly on both fronts. This aligns with previous research showing that liberals tend to emphasize moral values such as care and fairness, which support empathetic persuasion, while conservatives emphasize order and loyalty, which may limit persuasive flexibility Graham et al. (2009). Furthermore, conservatives’ higher self-confidence in judgment does not necessarily translate into effective interactive persuasion Ruisch and Stern (2020).